OpenStudy (anonymous):

two angles are supplementary if the sum of their measures is 180 degrees. the measure of the first angle is 54 degrees more than 3 times the second angle. find the measure of the first angle.

hero (hero):

1st angle = 3x + 54 2nd Angle = x 1st angle + 2nd Angle = 180 3x + 54 + x = 180 4x + 54 = 180

OpenStudy (anonymous):

a 4.5degrees b.148.5degrees c31.5 degrees d.175.5 degrees

OpenStudy (anonymous):

4x + 54 = 180 4x = 180 - 54 x = 126/4 x=31.5 so c 31.5 degrees

hero (hero):

Says find the measure of the 1st angle. You found the measure of the 2nd angle.

OpenStudy (anonymous):

so 31.5 times 3

OpenStudy (anonymous):

so first angle will be 3*31.5+54 = 94.5 + 54 =148.5 so ans is b.

OpenStudy (anonymous):

thank you so much i now understand how to do the problem a little more

hero (hero):

Next time draw a picture 

OpenStudy (anonymous):

ok will try that